Lulabalula:

Math- What is 75% of 200? 50 100 150 170

Lulabalula:

2.6 is what i got

Lulabalula:

2.6

Imagine:

50% of 200 =100

Lulabalula:

mhm

Imagine:

So what do you think the answer is.

Imagine:

So, \[ (75*200):100 =\]

Imagine:

Multiply what is in the () first.

Lulabalula:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @687vhlovdjfl \(\color{#0cbb34}{\text{Originally Posted by}}\) @Lulabalula \(\color{#0cbb34}{\text{Originally Posted by}}\) @687vhlovdjfl 150 \(\color{#0cbb34}{\text{End of Quote}}\) I wanted help not the exact answer I need to understand how to do it \(\color{#0cbb34}{\text{End of Quote}}\) 75 percent *200 = (75:100)*200 = (75*200):100 = 15000:100 = 150 \(\color{#0cbb34}{\text{End of Quote}}\) I know but I wanted help, not the exact answer @Imagine was helping I didn't need the answer bc he/she was helping me find it on my own

Lulabalula:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Imagine Multiply what is in the () first. \(\color{#0cbb34}{\text{End of Quote}}\) 15000

Imagine:

Multiply what is in the () first. you should get 1500 not 15000

Imagine:

wait my bad miss type.

Imagine:

75*20

Lulabalula:

ok

Lulabalula:

yeah 1,500

Imagine:

Then divide it by 100

Lulabalula:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Imagine Then divide it by 100 \(\color{#0cbb34}{\text{End of Quote}}\) 15

Imagine:

1500 divded by 10

Imagine:

These types are horrible

Imagine:

smh

Lulabalula:

ok 150

Imagine:

There's your answer ;)

Lulabalula:

oh wow thx that helped way more then the exact answer

badgirlhope:

@Lulabalula awesome figuring out the answer your self with help im proud of you

Lulabalula:

now i understand what to do

Imagine:

\[(75*20)\div10=150\]

Imagine:

Indeed.

Lulabalula:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Imagine \[(75*20)\div10=150\] \(\color{#0cbb34}{\text{End of Quote}}\) thanks so much

Imagine:

No problem :)